OpenStudy (anonymous):

pl help lim x-0 (integral from 0 to x cos(t^2) dt)/x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}[\int\limits_{0}^{x}\cos (t ^{2})dt]/x\]

OpenStudy (anonymous):

i think applying the l hospital rule will help but i cant differentiate the the upper part of the question that is\[d/dt[\int\limits_{0}^{x}\cos(t^2)dt]\]

OpenStudy (agent47):

Isn't that just the fundamental theorem of calculus? Derivative and integral cancel each other out, so you end up with Cos(x^2)

OpenStudy (anonymous):

but then limit would become cos(x^2)/x=1/0=not defined the answer wont match

OpenStudy (anonymous):

hey thanks i got it we have to differentiate both in numerator and denominator so on applying it would become cos(x^2)/1=1 thanks a lot

OpenStudy (agent47):

yes

OpenStudy (agent47):

I hadn't noticed that u didn't differentiate the denominator.

OpenStudy (anonymous):

yup same here i forgot to differentiate it